Homework #4 Due Friday, August 1, 2003

Least Squares

Answer all questions. Submit any pertinent hand calculations and/or Matlab code in a neat manner in order to help us give you partial credit. Readings: Van Loan 7.0-7.3, in section 7.3 skip the intro and start two sentences before Theorem 7.

1) Do problem P7.2.1 in the text. No programming necessary.

2) Do problem P7.2.3 in the text.

3) Do problem P7.3.3 in the text.

4) Do problem P7.3.4 in the text.

5) Do problem P7.3.5 in the text.

6) Do problem P7.3.6 in the text.